OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 3i Answer : 1 - 9i correct??

OpenStudy (anonymous):

multipy that and 1-3i

OpenStudy (anonymous):

so yu will get 10

OpenStudy (anonymous):

was gonna say that :P

OpenStudy (anonymous):

whatever katie... lies.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome. anytime.

OpenStudy (anonymous):

i was typyin was i not :P

jimthompson5910 (jim_thompson5910):

In general (a+b)(a-b) = a^2-b^2 So (1+3i)(1-3i) 1^2 - (3i)^2 1 - 9i^2 1 - 9(-1) 1+9 10 which means (1+3i)(1-3i) = 10

OpenStudy (anonymous):

thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome